In the emission line spectra of hydrogen atom how many lines can be accounted for by all possible electron transitions between five lowest energy levels within the atom.

Grade:11

1 Answers

Ravneet
5 years ago
 
Possible spectral lines= (n2-n1)(n2-n1+1)/2 this is the possible numbers. electron when jumps from n2=5 to n1=1 then according to the formula possible lines=(5-1)(5-1+1)/2=10
So the possibilities are 5 to 4 , 5 to 3 , 5 to 2, 5 to 1,  4 to 3 , 4 to 2 4 to 1, 3 to 2 , 3 to 1,  2 to 1